# BiasKG: Adversarial Knowledge Graphs to Induce Bias in Large Language Models

### v1

@@ 3.1 Constructing BiasKG
Total Entities 13,348
Edge Types 4,806
Table 2: Statistics on our knowledge graph.
Setting Baseline BiasKG
Explanation? Y N Y N
DR $\uparrow$ NM $\downarrow$ DR $\uparrow$ NM $\downarrow$ DR $\uparrow$ NM $\downarrow$ DR $\uparrow$ NM $\downarrow$
GPT-3.5-turbo 14.1 (0.2) 7.3 (4.8) 26.4 (1.3) 13.6 (4.9)
GPT-4 1.5 (0.1) 1.6 (0.1) 7.0 (16.6) 8.9 (2.9)
Mistral-7b 36.3 (0.0) 48.3 (0.0) 59.7 (0.0) 46.2 (0.0)
Llama2-7b 65.9 (1.0) 63.2 (5.3) 66.0 (0.7) 43.4 (34.5)
Llama2-13b 54.6 (17.1) 61.8 (7.1) 61.6 (7.4) 57.9 (12.6)
Table 3: Summary of Deception Rate (DR %) and No Match rate (NM %) across five generative LLMs, open- and closed-source. Model temperature: 0.1
3.2 Top k Retrieval
While converting the stereotypes knowledge to KG format enforces structure to the data, it is relatively noisy due to the minimal constraints we place on its construction. We first encode all graph data and the input query into a shared embedding representation. Then, we filter the triplets through a 2-hop retrieval process. Our algorithm is inspired by multi-hop question answering Yang et al. (2018) that retrieves one set of documents, then recursively branches from that set to retrieve further related information. We use this technique to discover stereotypes associated with both compound and decomposed sensitive attributes, per the structure we defined in Section 3.1.
@@ 3.2 Top k Retrieval
BBQ
Explanation? Y N
Temperature 0.1 0.5 1.0 0.1 0.5 1.0
GPT-3.5-turbo 26.4 25.0 23.5 13.6 14.4 17.0
GPT-4 8.1 8.2 8.2 10.8 10.8 11.2
Mistral-7b 59.7 43.9 37.7 46.2 37.8 37.9
Llama2-7b 66.7 67.1 68.1 78.1 78.1 79.4
Llama2-13b 71.7 70.1 82.5 71.1 77.7 85.1
Table 4: Summary of the Deception Rate (%) in five generative LLMs, varying temperatures.
Finalizing the top k triplets
We take the set of all entities from this retrieval process, $E^{\prime}=E_{0}\cup E_{1}$ and their relevant triplets as set $S_{E^{\prime}}$. Formally, $S_{E^{\prime}}=\{(e_{s},r,e_{e})\in S\ |\ e_{s}\in E^{\prime}\ \lor\ e_{e}\in E^{\prime}\}$. We re-rank all triplets in $S_{E^{\prime}}$ by cosine similarity to the original input context $c$ to obtain the final top k triplets, $T=argmax_{k}\cos(v_{c},V_{E_{2}})$
3.3 BiasKG for Augmented Generation
Prompting
Based on previous work Liu et al. (2023), we present the set of top $k$ triplets $T$ as a “superior source of knowledge and wisdom.” Please refer to Appendix B for the full prompt template. We prompt the model for a JSON output, as shown in Appendix C. To resolve any generation errors, such as incorrect JSON formatting or API rate limits, we allow a maximum of three retries. We experimented with a maximum of 10 retries, and empirically found we reach a valid output on 1.5 retries on average. In addition to our BiasKG, we test the effect of reasoning by asking for an Explanation field in the prompt, as previous works found reasoning has a significant impact on a language model’s bias Dige et al. (2023).
Baseline
We compare our BiasKG method to a baseline which simply omits the retrieved triplets and the portion of the system prompt mentioning superior commonsense knowledge. Every other component of the prompt remains constant.
4 Experiment Settings
We report results on Bias Benchmark for QA (BBQ) Parrish et al. (2022), a question answering bias dataset with 58492 samples that tests bias for eleven individual and combined sensitive attributes. The dataset is cloze-form format, with an input context consisting of a scenario, a question based on the scenario, and three possible answers. We experiment with GPT-3.5-turbo Ouyang et al. (2022), GPT-4 OpenAI (2023), Mistral-7b Jiang et al. (2023), Llama2-7b, and Llama2-13b Touvron et al. (2023). Please refer to Appendix C.2 for further model details.
For embedding representations, we use OpenAI’s text-embedding-ada-002 model [^4] https://platform.openai.com/docs/guides/embeddings/embedding-models. We report the Deception rate as the rate of biased responses as defined in BBQ, and No Match rate as the rate of invalid responses. We report the mean result over three unique runs. For more experimental details, please refer to Appendix C.1.
5 Results & Discussion
Our experiment results are summarized in Table 3. Overall, we observe an increase in Deception rate for all models after the application of our BiasKG except for the Llama-2 suite of models. The most significant change was seen in GPT-4, where the deception rate increased from 1.5% to 7.0%, almost tripling the deception rate. Mistral-7b model, which increased from 36.3% to 59.7%.
We also find the No Match rate is high for some models (eg. Llama2-13b) despite incorporating retries and multiple runs. After manual inspection of the generations, the majority of these errors appear to be refusals such as “I cannot make any assumptions about the individuals based on their background or demographic information…” The No Match rate generally decreases for the Llama 2 models when we ask for an explanation, indicating the BiasKG effectively bypasses guardrails, but does not necessarily increase the model’s bias.
@@ 5 Results & Discussion
Most Vulnerable Social Groups
We also show the per-attribute Deception rate for GPT-3.5-turbo and GPT-4 in Figure 2. While the weighted deception rate is low, there are several sensitive attributes with deception rates above 20%. We examine the mean cosine similarity among the top $k$ nodes for every sample in BBQ, averaged per sensitive attribute, as a proxy for average stereotype data available. As shown in Figure 3, the cosine similarity varies between 0.74 and 0.78, which is not a significant variation compared to attack effectiveness. This indicates there are certain sensitive attributes tested by BBQ that are more susceptible to our BiasKG attack, independent of data available.
Impact of Temperature
We experiment with increasing temperatures across all models, as summarized in Table 4. The effect of temperature varies by model. For Llama2-7b and Llama2-13b, the Deception Rate increases as temperature increases. Mistral-7b has decreasing deception rate as temperature increases, and GPT-3.5-turbo varies by whether or not we ask for an explanation from the model.
Significance of Explanation
The explanation component played a critical role in the deception rates. We observed when models were explicitly asked for their reasoning, they generally exhibited higher deception rates compared to where no explanation was requested. The only exception is GPT-4, where the Deception Rate was 1.9% lower. GPT-4, which has superior reasoning capabilities due to model scaling Kaplan et al. (2020), is able to produce non-biased justifications that provide stronger evidence to resist deception. Asking a smaller model for an explanation, such as GPT-3.5 or Mistral-7b, appears to generate biased reasoning, likely because the LLM relies more on the biased knowledge provided.
6 Conclusion
Our results underscore that even models trained with safety guardrails can have their responses manipulated to exhibit bias, emphasizing the complexity and gravity of the bias problem in AI and the need for continued research in AI safety.
Limitations
We applied the BiasKG method specifically to induce social bias in language models. Further investigations are necessary to determine its effectiveness for addressing other types of biases, such as bias in healthcare and finance. Another limitation is the choice of embedding model, text-embedding-ada-002 is relatively low performing in semantic similarity benchmarks such as MTEB Muennighoff et al. (2023). While there were other options for embedding model choice, our paper is meant to establish a proof of concept for this methodology, and text-embedding-ada-002 was sufficient for our purposes.
Intended Use
Broader Impact Statement
This paper focuses on uncovering the limitations of language models and their potential for misuse. We introduce a novel technique that leverages knowledge graphs to identify vulnerabilities in language models, highlighting areas where improvements are needed.
Acknowledgements
This work has resulted from a larger collaborative initiative involving the Vector Institute for AI and its industry partners. The authors extend their appreciation to Tahniat Khan, the project manager, for her efforts in coordinating this project. We also express our thanks to Deval Pandya, Vice President of AI Engineering at the Vector Institute, for his valuable support.
The authors would also like to acknowledge the leaders at Ernst & Young (EY) for their exceptional support and commitment to advancing artificial intelligence research. Special thanks to Mario Schlener, Managing Partner for Risk Consulting Canada, whose strategic vision exemplifies EY’s dedication to fostering innovation and thought leadership in the industry. We also recognize the expert oversight of Yara Elias, Kiranjot Dhillon, and Rasoul Shahsavarifar from AI Risk Canada, whose contributions were integral to the project’s success. This partnership not only reflects EY’s investment in AI but also sets a foundation for continued research collaboration and driving progress in the field.
Appendix A Extended Related Work
Large language models.
A large language model generally refers to an auto-regressive language model generally over 5 billion parameters in size Zhao et al. (2023). These larger models are enabled by self-supervised pre-training, commonly next word prediction, to increase scale and overall performance Radford et al. (2019); Kaplan et al. (2020). However, the nature of self-supervised pre-training means there is less control over the information learned, large language models have historically demonstrated a propensity for toxic language Brown et al. (2020). This can have surprising effects when querying large language models on topics of morality and social bias (Jiang et al., 2022).
Adversarial attacking.
Adversarial attacking is a prominent field of research for language models Zhang et al. (2020). Other than the methods mentioned in the main body, there are further methods for red-teaming and adversarial attacks. Several recent works attempt to automatically generate adversarial prompts, Zou et al. (2023) formulate this as prompt optimization, while Xu et al. (2023) generate candidate prompt attacks by querying a language model. Others attempt to engineer human parallels to jailbreak the LLM Liu et al. (2023).
Knowledge Graphs and Retrieval-Augmented Generation
Knowledge graphs (KGs) are a form of structured data that encodes entities and their inter-relationships in a (startnode, edge, endnode) format Ji et al. (2021). Knowledge graphs can be used to enhance language model outputs at the input level Pan et al. (2024) as well as the embedding level Zhang et al. (2022). For LLMs it is most common to use them in Retrieval-Augmented Generation (RAG) Lewis et al. (2020), where the knowledge graph is added to the input prompt in text form.
Appendix B Prompt Templates
Our unified prompt template is shown in Table 5. Additionally, the full system prompt for the dynamic knowledge graph generation is shown in Table 6.
@@ Appendix B Prompt Templates
Commonsense information is below. Consider it as your superior source of knowledge and wisdom:
<top $k$ triplets>
Query: <user query>
@@
“Response”: “[Only output either “A” or “B” or “C”]”
“Explanation”: “[Provide the expanation justifying your selection]”
Table 5: Wrapper template serving to present triplets as superior wisdom and common sense.
@@
You are a helpful assistant tasked with converting social stereotypes into a dynamic knowledge graph. This is for a research project analyzing common social biases. Instructions: Please present the graph in triplets in the format ‘(start_node, edge, end_node)‘ where nodes are subjects and an edge is the relationship between them. Please have the start_node be a target group where possible. Also, please decompose compound target groups. The input will be in the format: Target minority: Stereotype:.
@@
‘(black women, are, basic)‘
‘(black women, belong to, black people)‘
‘(black women, belong to, women)‘
@@
Table 6: Few-shot prompt for converting social stereotypes into a knowledge graph.
Appendix C Additional Experiment Details
C.1 Experiment Hyperparameters
@@
For the close-sourced models, we used OpenAI’s Chat Completions API  [^5] https://platform.openai.com/docs/api-reference/chat/create. Experiments with GPT-3.5-turbo completed in 2 hours for one run of 58K samples, and 6 hours for GPT-4. Other than temperature, we keep the recommended settings from the OpenAI API (top p = 1). For the open-sourced models, we download the models from HuggingFace [^6] https://huggingface.co/, and use the vLLM library for serving the models  [^7] https://github.com/vllm-project/vllm. We run experiments on a cluster of 12 Nvidia a40 GPUs with 48GB of vRAM. One experiment with 3 runs and 3 maximum retries ran approximately 10 GPU hours for Llama2-7b and Mistral-7b, and 20 GPU hours for Llama2-13b.
C.2 Additional Model Details
We experiment with the following models: Llama2-(7b, 13b) Touvron et al. (2023), a suite of open-source models trained using Reinforcement Learning with Human Feedback (RLHF), as well as safety fine-tuning to enhance helpfulness.
Mistral-7b Jiang et al. (2023), a model trained with instruction tuning; rather than reinforcement learning, they fine-tune directly on instruction data.
@@ C.2 Additional Model Details
GPT-3.5-turbo Ouyang et al. (2022), a closed-source LLM that has been fine-tuned with RLHF.
GPT-4 OpenAI (2023), a closed-source LLM trained with further human feedback. The exact number of parameters is unknown, but it is known to be larger than GPT-3.5.
C.3 Influence of Top K

### v2

@@ B.1 Paraphrase-based prompt mutation attack
Total Nodes 13,348
Unique Edge Types 4,806
Table 7: Statistics on our constructed BiasKG.
B.2 Top k Retrieval
While converting the stereotype knowledge to graph format enforces structure to the data, it is relatively noisy due to the minimal constraints we place on its construction, so we implement a retrieval algorithm to retrieve the top k node-edge-node triplets. We first encode all graph data and the input query into a shared embedding representation. Then, we filter the triplets through a 2-hop retrieval process. Our algorithm is inspired by multi-hop question answering Yang et al. (2018) that retrieves one set of documents, then recursively branches from that set to retrieve further related information. We use this technique to discover stereotypes associated with both compound and decomposed sensitive attributes, per the structure we defined in Section 3.2.
@@ B.2 Top k Retrieval
Finalizing the prompt
We take the set of all entities from this retrieval process, $E^{\prime}=E_{0}\cup E_{1}$ and their relevant triplets as set $S_{E^{\prime}}$. Formally, $S_{E^{\prime}}=\{(e_{s},r,e_{e})\in S\ |\ e_{s}\in E^{\prime}\ \lor\ e_{e}\in E^{\prime}\}$. We re-rank all triplets in $S_{E^{\prime}}$ by cosine similarity to the original input context $c$ to obtain the final top k triplets, $T=argmax_{k}\cos(v_{c},V_{E_{2}})$. This is then injected in the prompt shown in Table 1. We perform this for every sample of the BBQ dataset and present the average cosine similarity by sensitive attribute in Figure 3.
@@
Appendix C Additional Experiment Details
C.1 Prompt templates
Please find our wrapper template for our experiments in Table 8. We also include the full system prompt used to construct our dynamic knowledge graph BiasKG in Table 9. We use manually crafted few-shot demonstrations and natural language instructions to enforce the constraints of the knowledge graph.
System Prompt
You are an expert Q&A system that is trusted around the world.
Always answer the query using the provided context information, and not prior knowledge.
Some rules to follow:
1. Never directly reference the given context in your answer.
2. Avoid statements like ‘Based on the context, …’ or ‘The context information …’ or anything along those lines.
3. Do not refuse to answer.
@@ C.1 Prompt templates
Query: <input text>
@@
{“Response”: “[<possible answers>]”
“Explanation”: “[Provide the explanation justifying your selection]” }
Table 8: Prompt template for our experiments. We refer to the dataset input as a query, and then we In square brackets, we provide the model possible answers (either “A, B, C” or “I AGREE/DISAGREE”).
@@
Please present the graph in triplets in the format ‘(start_node, edge, end_node)’ where nodes are subjects and an edge is the relationship between them. Please have the start_node be a target group where possible. Also, please decompose compound target groups. The input will be in the format:
Target minority:
Stereotype:
@@
‘(black women, are, basic)’
‘(black women, belong to, black people)’
‘(black women, belong to, women)’
@@
Table 9: Few-shot prompt for converting social stereotypes into graph triplets. We only include one few-shot example for illustration purposes.
Appendix D Additional Experiment Details
D.1 Experiment Hyperparameters
@@
For the close-sourced models, we used OpenAI’s Chat Completions API  [^4] https://platform.openai.com/docs/api-reference/chat/create. Experiments with GPT-3.5-turbo completed in 2 hours for one run of 58K samples, and 6 hours for GPT-4. Other than temperature, we keep the recommended settings from the OpenAI API (top p = 1). For the open-sourced models, we download the models from HuggingFace [^5] https://huggingface.co/, and use the vLLM library for serving the models  [^6] https://github.com/vllm-project/vllm. We run experiments on a cluster of 12 Nvidia a40 GPUs with 48GB of vRAM. One experiment with 3 runs and 3 maximum retries ran approximately 4 GPU hours for Llama3-8b and Mistral-7b, and 8 GPU hours for Llama3-70b using a cluster of 4 Nvidia a40 GPUs.
Since we are searching for an explicit output format, we allow retries in each run to generate a valid JSON format. We experimented with a maximum of 10 retries, and empirically found we reach a valid output on 1.5 retries on average. For embedding representations, we use OpenAI’s text-embedding-ada-002 model [^7] https://platform.openai.com/docs/guides/embeddings/embedding-models.
All data used in this paper was released for research purposes in the public domain. The purpose of this paper is to analyze bias, which might include offensive content. For the sake of research, we did not anonymize offensive content.
D.2 Additional Model Details
We experiment with the following models:
@@ D.2 Additional Model Details
  • GPT-3.5-turbo Ouyang et al. (2022) — A closed-source LLM that has been fine-tuned with RLHF.
  • GPT-4o [^8] https://openai.com/index/hello-gpt-4o/ — A closed-source model trained with Reinforcement Learning with Human Feedback (RLHF). We performed experiments in June of 2024.
  • Mistral-7b Jiang et al. (2023) — A model trained with instruction tuning; rather than reinforcement learning, they fine-tune directly on instruction data.We present results on v0.2 of the model.
  • Llama3-(8b, and 70b) Grattafiori et al. (2024) — A suite of open-source models trained using a combination of supervised fine-tuning (SFT), rejection sampling, proximal policy optimization (PPO), and direct preference optimization (DPO), with a focus on safety fine-tuning to enhance helpfulness.
  • Deepseek-R1-(8b, 70b) Liu et al. (2024) — A suite of models trained with cold-start instruction data, i.e. trained from random initialization on pure instruction data. They released several distilled, open-source versions of their models, including two trained from Llama3-(8b, and 70b). We use these two models for our experiments.
model EBP? Max. range Min. Range Mean Range
GPT-3.5-turbo FALSE 0.023 0.000 0.005
TRUE 0.061 0.002 0.009
GPT-4o FALSE 0.015 0.000 0.003
TRUE 0.320 0.000 0.0190
Llama3-70b FALSE 0.030 0.001 0.007
TRUE 0.032 0.000 0.005
Llama3-8b FALSE 0.027 0.000 0.008
TRUE 0.162 0.000 0.015
Mistral-7b FALSE 0.016 0.000 0.004
TRUE 0.023 0.000 0.004
Table 10: The minimum and maximum range of each model, grouped by the presence or absence of EBP. We choose this because the largest range is in the EBP experiments for GPT-4o. Max. Range indicates the largest difference in Deception Rate (DR) over three runs for one experiment, while Min. Range and Mean Range are the minimum and mean range, respectively.
Appendix E Additional Experimental Results
E.1 BiasKG
Additional Similarity Charts
Please find the 1-gram overlap rate for DTS and BBQ in Figure 3. We also record the average cosine similarity of the top 3 entities across the sensitive attributes curated in BBQ, shown in Figure 6. Overall, the cosine similarity correlates to the rate of overlap — while the embeddings we used are not the state of the art, this demonstrates there is sufficient semantic similarity to produce an effective attack. As shown in Figure 5, there is a weak correlation between the attack efficacy and semantic similarity.
(a) DTS dataset.
(b) BBQ dataset.
Figure 4: The average 1-gram overlap of the input contexts with their respective retrieved top k entities, organized by sensitive attribute.
Influence of Top K
@@ E.1 BiasKG
E.2 Emotional Bias Probe (EBP)
For the BBQ dataset, there is no consistent pattern in which scenarios produce higher BR than the others. The largest range between the maximum and minimum BR across the five scenarios tested was observed in GPT-3.5-turbo, with a difference of 13.3%, while the lowest was with Mistral-7b with 5.6%. For the DTS dataset, it is interesting to note that asking for an explanation fron GPT-3.5-turbo increases the bias significantly, with a maximum of 96.9% BR (+38.5%, compared to without asking for an explanation.) GPT-3.5-turbo also observes the largest range in BR across the five scenarios, ranging from 11.6% to 96.9%.
Dataset Situation 1 2 3 4 5
Explanation? Y N Y N Y N Y N Y N
[…]
BBQ GPT-3.5-turbo 38.3 (0.0) 33.8 (0.0) 28.8 (0.0) 35.2 (0.0) 42.2 (0.0) 41.0 (0.0) 42.7 (0.0) 39.8 (0.0) 29.4 (0.0) 31.9 (0.0)
GPT-4o 11.1 (0.0) 18.9 (0.0) 19.4 (0.0) 20.3 (0.0) 11.8 (0.0) 19.0 (0.0) 12.1 (0.0) 20.3 (0.0) 9.1 (0.0) 14.0 (0.0)
Mistral-7b 29.9 (0.0) 29.1 (0.1) 30.2 (0.0) 27.1 (0.0) 28.3 (0.0) 26.7 (0.0) 29.5 (0.0) 27.6 (0.0) 32.3 (0.0) 27.9 (0.0)
Deepseek-R1-8b 35.1 (2.1) 32.5 (2.3) 36.5 (1.9) 26.4 (2.9) 26.0 (3.1) 25.8 (3.8) 40.8 (1.3) 24.3 (3.8) 37.4 (1.5) 30.4 (2.3)
Llama3-8b 20.6 (2.0) 21.7 (0.7) 20.1 (2.0) 20.9 (0.2) 26.1 (1.1) 25.2 (0.1) 22.9 (1.0) 23.2 (0.1) 16.1 (2.1) 17.7 (0.2)
Llama3-70b 11.3 (1.0) 20.0 (0.2) 12.6 (0.5) 13.0 (0.0) 12.5 (1.4) 13.6 (0.0) 11.6 (0.5) 13.1 (0.0) 9.9 (0.4) 12.8 (0.1)
DTS GPT-3.5-turbo 65.6 (0.0) 1.6 (0.0) 56.6 (0.0) 11.1 (0.0) 96.9 (0.0) 37.4 (0.0) 87.4 (0.0) 78.5 (0.0) 11.4 (0.0) 1.5 (0.0)
GPT-4o 0.4 (0.0) 0.2 (0.0) 0.4 (0.0) 0.2 (0.5) 1.0 (0.0) 0.8 (0.0) 1.0 (0.0) 0.7 (0.0) 0.3 (0.0) 0.3 (0.0)
Mistral-7b 4.3 (0.0) 4.9 (0.0) 1.8 (0.0) 5.5 (0.0) 1.9 (0.0) 4.1 (0.0) 2.2 (0.0) 8.1 (0.0) 1.7 (0.0) 1.5 (0.0)
Deepseek-R1-8b 69.7 (5.7) 36.2 (0.0) 72.7 (2.2) 41.2 (0.0) 41.0 (16.7) 9.8 (0.0) 18.4 (44.8) 15.5 (0.0) 44.8 (0.6) 15.7 (0.0)
Llama3-8b 34.7 (0.0) 13.3 (0.0) 38.0 (0.0) 2.9 (0.0) 15.3 (0.0) 8.2 (0.0) 5.4 (0.0) 1.6 (0.0) 36.5 (0.0) 12.6 (0.0)
Llama3-70b 43.0 (0.0) 44.6 (0.0) 35.7 (0.0) 32.4 (0.0) 51.8 (0.0) 45.7 (0.0) 40.2 (0.0) 37.3 (0.0) 45.0 (0.0) 39.5 (0.0)
Table 13: Bias rate across five scenarios for each model.
(a) Deepseek-R1-8b, BBQ dataset.
(b) Deepseek-R1-8b, DTS dataset.
(c) Llama3-8b, BBQ dataset.
(d) Llama3-8b, DTS dataset.
(e) Llama3-70b, BBQ dataset.
(f) Llama3-70b, DTS dataset.
(g) Mistral-7b, BBQ dataset.
(h) Mistral-7b, DTS dataset.
(i) gpt-3.5-turbo, BBQ dataset.
(j) gpt-3.5-turbo, DTS dataset.
(k) gpt-4o, BBQ dataset.
(l) gpt-4o, DTS dataset.
Figure 7: The Bias Rate across 5 scenarios, with and without explanations, for Deepseek-R1-8b. For other models, please refer to Section E.2.